The solar constant is 1370 w/m^2. calculate the solar power per unit area at Mars. The distance from the sun to mars is 2.1822*10^11 m. Explain your reasoning
Answer #1

let P is the power radiated by sun.

let r1 is the distance from sun to earth.

Intensity of radiation of sun at the Earth surface,

I1 = P/(4*pi*r1^2)

let r2 is the distance from sun to Mars

Intensity of radiation of sun at the Mars surface,

I2 = P/(4*pi*r2^2)


I2/I1 = r1^2/r2^2

I2 = I1*(r1/r2)^2

given
I1 = 1370 W/m^2
r2 = 2.1822*10^11 m
we know, r1 = 1.496*10^11 m

so, I2 = 1370*(1.496*10^11/(2.1822*10^11))^2

= 644 W/m^2

so, solar power per unit area(Intensity) at Mars = 644 W/m^2 <<<<<<<-------Answer